Suppose that 20% of people have a dog, 29% of people have a cat, and 13% own both. What is the probability that someone owns a dog or a cat?
Answer #1

P(Dog) = 0.2

P(Cat) = 0.29

P(Dog and Cat) = 0.13

Probability that someone owns a dog or a cat = P(Dog or Cat) = P(Dog) + P(Cat) - P(Dog and Cat)

= 0.2 + 0.29 - 0.13

= 0.36